OpenStudy (anonymous):

PLEASE HELP ! Simplify the expression sin^2x-1/cos(-x) answers : -sin x cos x sin x -cos x

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ sin^2(\theta)+cos^2(\theta)=1\implies sin^2(\theta)-1=-cos^2(\theta)\\ \quad \\ cos(-x)=cos(x)}}\\ \quad \\ \quad \\ \cfrac{sin^2(x)-1}{cos(-x)}\) what do you think?

OpenStudy (anonymous):

i have no clue?

OpenStudy (jdoe0001):

well, why not use those identities? the pythagorean one and the symmetric one

OpenStudy (jdoe0001):

then simplify